OpenStudy (psymon):

Find a set of parametric equations for a line that passes through the point <-2,5,4> and is parallel to the xy-plane and yz-plane. Find a set of parametric equations for a line that passes through the point <-4,5,2> and is perpendicular to the plane given by -x+2y+z = 5 I wanna make sure I'm doing these correctly, not enough examples.

OpenStudy (tkhunny):

Parallel to xy and yz? So, only y can be altered?

OpenStudy (dumbcow):

^^ correct so the line would be: x(t) = -2 y(t) = t+5 z(t) = 4 part2) vectors which are perpendicular have dot product of 0 plane vector is coefficients --> <-1,2,1> line vector is <a,b,c> -a +2b +c = 0 a = 1 b=1 c = -1 the point (-4,5,2) is initial value when t=0 x(t) = t -4 y(t) = t+5 z(t) = 2 - t